OpenStudy (anonymous):

25 + (−6) + (−5) = 25 + (−5) + c

OpenStudy (anonymous):

c= -6

OpenStudy (anonymous):

:)

OpenStudy (anonymous):

ok thx ur so helpful :)

OpenStudy (anonymous):

No problem any time :)

OpenStudy (anonymous):

I have two more questions

OpenStudy (anonymous):

−(7 + y) = −7 + (−3)

OpenStudy (rushwr):

3

OpenStudy (anonymous):

ok thx

OpenStudy (anonymous):

7 + x + (−17) = 2

OpenStudy (anonymous):

y=3 :)

OpenStudy (rushwr):

\[-(7+y) = -7 + (-3) \] \[-7-y = -10\] \[y= 10 -7 = 3\]

OpenStudy (anonymous):

here is the last question

OpenStudy (anonymous):

7 + x + (−17) = 2

OpenStudy (anonymous):

x=12

OpenStudy (anonymous):

:)

OpenStudy (anonymous):

thx :)

OpenStudy (anonymous):

Anytime :)